Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used iGitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used iRequired: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overh$ $ 135,000 81,000 Purcnase of raw materiais Direct labor cost Manufacturing overhead costs: Indirect labor Property taxes DeInsurance Rent, building $ $ 7,200 31,000 points Raw Materials Work in Process Finished Goods Beginning Ending $ 27,000 $ 18,
Answer #1

1) Predetermine overhead rate = Manufacturing overhead/Direct material used = 102000/85000 = 120% of direct material used

2) Direct material used = 27000+135000-18000 = 144000

Applied overhead = 144000*1.2 = 172800

Actual overhead = 177300

Under applied overhead = 4500

3) Schedule of cost of goods manufactured

Direct material
Beginning raw material 27000
Raw material purchase 135000
Raw material available for use 162000
Less; Ending raw material -18000
Direct material used 144000
Direct labor 81000
Manufacturing overhead 172800
Total manufacturing cost 397800
Beginning work in process 46000
Total Cost of work in process 443800
Less: Ending work in process -40000
Cost of goods manufactured 403800

4) Schedule of cost of goods sold

Beginning finished goods 73000
Cost of goods manufactured 403800
Cost of goods available for sale 476800
Less: Ending finished goods -60000
Unadjusted cost of goods sold 416800

5) Calculate following

Direct material 8200
Direct labor 21960
Manufacturing overhead (8200*1.2) 9840
Work in process inventory 40000